# Jian-hao Kanga, Zhun Goub and Nan-jing Huangc

### v1

@@ 4 Equilibrium strategies to the optimization problem
In this section, we first attempt to address the delayed reinsurance and investment optimization problem with random risk aversion in a general utility function framework as far as possible. Subsequently, in order to obtain analytical equilibrium reinsurance and investment strategies, we narrow our focus to analyze the exponential utility function and the power utility function, respectively.
@@
Having derived the pseudo HJB equation (3.13) that characterizes the equilibrium reinsurance and investment strategy, we can perform partial differentiation on the terms within the brackets of equation (3.13) with respect to $q$ and $\pi$ separately. This step enables us to obtain
@@
$\hat{\pi}(t)=-\frac{(\mu-r)U_{x}(t,x,m_{1})}{\sigma^{2}x\left[U_{xx}(t,x,m_{1})-\int(\iota^{\gamma})^{{}^{\prime}}(Y^{\gamma}(t,x,m_{1}))(Y_{x}^{\gamma}(t,x,m_{1}))^{2}\mathrm{d}\Gamma(\gamma)\right]},$ (4.15)
where we denote by $(\hat{q}(t),\hat{\pi}(t)):=(\hat{q}(t,x,m_{1}),\hat{\pi}(t,x,m_{1}))$ the candidate equilibrium reinsurance and investment strategy. It follows from the equation (C) that
$U(t,x,m_{1})=H(t,x,m_{1})=\int(\varphi^{\gamma})^{-1}\left(Y^{\gamma}(t,x,m_{1})\right)\mathrm{d}\Gamma(\gamma).$
Then, through the system of equations (3.12), equations (4.14) and (4.15) admit the following reformulation
@@ 4.1 The exponential utility
$\varphi^{\gamma}(\cdot)=-\frac{1}{\gamma}e^{-\gamma(\cdot)},$ (4.18)
Consequently, equations (4.16) and (4.17) reduce to
@@
We postulate an ansatz where $Y^{\gamma}(t,x,m_{1})=-\frac{1}{\gamma}e^{g_{1}^{\gamma}(t)(x+\beta m_{1})+g_{2}^{\gamma}(t)}$ with $g_{1}^{\gamma}(t)$ and $g_{2}^{\gamma}(t)$ being deterministic functions of $t\in[0,T]$ with $g_{1}^{\gamma}(T)=-\gamma$ and $g_{2}^{\gamma}(T)=0$. It should be noted that the superscript $\gamma$ on functions $g_{1}^{\gamma}(t)$ and $g_{2}^{\gamma}(t)$ signifies their dependence on $\gamma$. This notation does not imply that $g_{1}(t)$ and $g_{2}(t)$ are raised to the $\gamma$-th power.
@@
$\left\{\begin{aligned} &Y^{\gamma}_{t}(t,x,m_{1})=-\frac{1}{\gamma}e^{g_{1}^{\gamma}(t)(x+\beta m_{1})+g_{2}^{\gamma}(t)}\left[\frac{\partial g_{1}^{\gamma}(t)}{\partial t}\left(x+\beta m_{1}\right)+\frac{\partial g_{2}^{\gamma}(t)}{\partial t}\right],\\ &Y^{\gamma}_{x}(t,x,m_{1})=-\frac{1}{\gamma}g_{1}^{\gamma}(t)e^{g_{1}^{\gamma}(t)(x+\beta m_{1})+g_{2}^{\gamma}(t)},\\ &Y^{\gamma}_{xx}(t,x,m_{1})=-\frac{1}{\gamma}(g_{1}^{\gamma}(t))^{2}e^{g_{1}^{\gamma}(t)(x+\beta m_{1})+g_{2}^{\gamma}(t)},\\ &Y^{\gamma}_{m_{1}}(t,x,m_{1})=-\frac{1}{\gamma}\beta g_{1}^{\gamma}(t)e^{g_{1}^{\gamma}(t)(x+\beta m_{1})+g_{2}^{\gamma}(t)}\end{aligned}\right$.
@@
By inserting the partial derivatives of $Y^{\gamma}(t,x,m_{1})$ into (3.10), we obtain that
@@
where $\hat{q}$ and $\hat{\pi}$ are provided by equations (4.19) and (4.20). We rewrite the above equation into the subsequent form
$0=$ $\frac{\partial g_{1}^{\gamma}(t)}{\partial t}(x+\beta m_{1})+\frac{\partial g_{2}^{\gamma}(t)}{\partial t}+\left[\hat{\pi}(\mu-r)x+a\eta+a\eta_{2}\hat{q}\right]g_{1}^{\gamma}(t)+0.5(b^{2}\hat{q}^{2}+\hat{\pi}^{2}x^{2}\sigma^{2})(g_{1}^{\gamma}(t))^{2}$
$+\left[(A+\beta)x+(B-\alpha\beta)m_{1}+(C-e^{-\alpha h}\beta)m_{2}\right]g_{1}^{\gamma}(t).$ (4.21)
Thus, under the following assumption
  • $(A2)$
    $C=\beta e^{-\alpha h}$ and $Be^{-\alpha h}=(\alpha+A+\beta)C$,
the equation (4.1) becomes
@@
Remark 4.1.
We would like to mention that the assumption $(A2)$ can be used to ensure the existence of a closed-form solution for the equation (3.10). Although such an assumption limits the generality to some extent, it is beneficial to transform the delayed reinsurance and investment optimization problem into a finite-dimensional one. Furthermore, it is a sufficient condition to make sure that $V$ depends solely on $(t,x,m_{1})$.
One can see that $\hat{q}$ and $\hat{\pi}x$ are independent of $x$ and $m_{1}$ due to equations (4.19) and (4.20). Thus, applying the coefficient comparison method to the terms containing $x+\beta m_{1}$ in (4.1) generates
@@
where $\mathbb{E}[\gamma]=\int\gamma\mathrm{d}\Gamma(\gamma)$ and
$D=\gamma\left[\frac{(\mu-r)^{2}}{\sigma^{2}\mathbb{E}[\gamma]}+\frac{a^{2}\eta^{2}_{2}}{b^{2}\mathbb{E}[\gamma]}\right]-0.5\gamma^{2}\left[\frac{(\mu-r)^{2}}{\sigma^{2}\mathbb{E}^{2}[\gamma]}+\frac{a^{2}\eta^{2}_{2}}{b^{2}\mathbb{E}^{2}[\gamma]}\right].$
By using the equation (C), one can further derive that the candidate equilibrium value function is
@@
In addition, provided the expression for the candidate equilibrium reinsurance and investment strategy, one can present the sensitivity analysis on the effects of the delay parameters on $(\hat{q}(t),\hat{\pi}(t)x)$, which are given as follows.
@@
$sign\left(\frac{\partial\hat{f}(t)}{\partial\alpha}\right)=\begin{cases}1,\quad\quad\text{for}\;\quad\,\alpha>-\frac{1}{h}\ln\frac{1}{h},\\ 0,\quad\quad\text{for}\;\quad\,\alpha=-\frac{1}{h}\ln\frac{1}{h},\\ -1,\quad\;\text{for}\;\quad\,\alpha<-\frac{1}{h}\ln\frac{1}{h},\end{cases}$ (4.28)
@@
$sign\left(\frac{\partial\hat{f}(t)}{\partial\beta}\right)=\begin{cases}1,\quad\quad\text{for}\;\quad\,h<-\frac{1}{\alpha}\ln(1-r-\alpha),\\ 0,\quad\quad\text{for}\;\quad\,h=-\frac{1}{\alpha}\ln(1-r-\alpha),\\ -1,\quad\;\text{for}\;\quad\,h>-\frac{1}{\alpha}\ln(1-r-\alpha).\end{cases}$ (4.29)
@@
Assume that the utility function $\varphi^{\gamma}$ is given by the equation (4.18) with $\gamma\in[\epsilon_{1},\infty)$ and the assumption $(A2)$ holds, where $\epsilon_{1}$ is a very small positive constant. Then the equilibrium reinsurance and investment strategy and the corresponding equilibrium value function are given by equations (4.25) and (4.27), respectively.
@@
Note that the equilibrium reinsurance strategy in (4.31) is consistent with the one in the equation (2.9) of [2] and the equilibrium investment amount $\hat{\pi}(t)x$ in (4.31) coincides with the ones with $\beta=0$ in equations (3.11) and (3.14) of [1], which means that under singular risk aversion, the equilibrium strategy solving the problem (3.5) is mathematically equivalent to the optimal strategy solving the problem (3.4).
@@ 4.2 The power utility
We propose an ansatz in which $Y^{\gamma}(t,x,m_{1})=\frac{1}{1-\gamma}(g^{\gamma}(t))^{\gamma}(x+\beta m_{1})^{1-\gamma}$ with $g^{\gamma}(t)$ being a deterministic function of $t\in[0,T]$ with $g^{\gamma}(T)=1$. The superscript $\gamma$ in the function $g^{\gamma}(t)$ denotes its reliance on the parameter $\gamma$. It’s crucial to emphasize that this notation doesn’t indicate that $g(t)$ is exponentiated to the power of $\gamma$.
@@
$\left\{\begin{aligned} &Y^{\gamma}_{t}(t,x,m_{1})=\frac{\gamma}{1-\gamma}(g^{\gamma}(t))^{\gamma-1}(x+\beta m_{1})^{1-\gamma}\frac{\partial g^{\gamma}(t)}{\partial t},\\ &Y^{\gamma}_{x}(t,x,m_{1})=(g^{\gamma}(t))^{\gamma}(x+\beta m_{1})^{-\gamma},\\ &Y^{\gamma}_{xx}(t,x,m_{1})=-\gamma(g^{\gamma}(t))^{\gamma}(x+\beta m_{1})^{-\gamma-1},\\ &Y^{\gamma}_{m_{1}}(t,x,m_{1})=\beta(g^{\gamma}(t))^{\gamma}(x+\beta m_{1})^{-\gamma}.\end{aligned}\right$.
When we substitute these partial derivatives into the candidate equilibrium reinsurance and investment strategy, it produces
@@
Inserting the above partial derivatives of $Y^{\gamma}(t,x,m_{1})$ into the equation (3.10) allows us to derive
@@
It is worth mentioning that equations (4.38) and (4.2) demonstrate that because of the linear architecture of the ODE for $g^{\gamma}(t)$, which is presented in the equation (4.36), we are able to transform the infinite-dimensional system of ODEs into a description using a single equation. Nevertheless, the equation (4.2) implicitly defines $\varpi(t,\Gamma)$ through a nonlinear operator equation. While closed-form solutions for $\varpi(t,\Gamma)$ may exist under special distributional assumptions on $\Gamma$, the general case resists analytical treatment. Therefore, we attempt to discretize $\Gamma$. For discrete distributions, the outer integral over $\Gamma$ reduces to a finite sum, transforming the system of ODEs into a more tractable form.
@@ 4.2.1 Special case with $n$ possible risk aversions
Suppose the random risk aversion is considered to conform to an $n$-point distribution having possible outcomes $\gamma_{i}>0\;(i=1,\cdot\cdot\cdot,n)$, which occur with probabilities $p_{i}=p(\gamma_{i})$ and $\sum\nolimits_{i=1}^{n}p_{i}=1$. Then, we are able to derive
@@
and the corresponding equilibrium value function is
$V(t,x,m_{1})=(x+\beta m_{1})\sum\limits_{i=1}^{n}(g^{\gamma_{i}}(t))^{\frac{\gamma_{i}}{1-\gamma_{i}}}p_{i},$
where $\varpi(t,\Gamma)$ is given by the equation (4.40).
@@
If $\alpha=h=B=C=\beta=0$, which indicates that there is no inclusion of any delay within the model, then the equilibrium reinsurance and investment strategy becomes
$\hat{q}(t)=\frac{a\eta_{2}x}{b^{2}\varpi(t,\Gamma)},\quad\hat{\pi}(t)=\frac{(\mu-r)}{\sigma^{2}\varpi(t,\Gamma)}$
and the corresponding equilibrium value function is characterized by
$V(t,x)=x\sum\limits_{i=1}^{n}(g^{\gamma_{i}}(t))^{\frac{\gamma_{i}}{1-\gamma_{i}}}p_{i},$
where $\varpi(t,\Gamma)$ is expressed by the equation (4.40) with $g^{\gamma_{i}}(t)$ satisfying
@@
It follows from equations (4.40) and (4.41) that the solution to $g^{\gamma_{i}}(t)$ is influenced by the model parameters of both the risky asset and the insurance market. As a result, $\varpi(t,\Gamma)$ and $(\hat{q}(t),\hat{\pi}(t))$ are also affected by these parameters. This implies that when the financial market and the insurance market are independent, the equilibrium investment strategy is influenced not only by the parameters of the risky asset but also by those of the insurance market. The same phenomenon exists for the equilibrium reinsurance strategy. However, these phenomena do not exist in the case of the exponential utility function with random risk aversion (see Theorem 4.1). Furthermore, we would like to mention that in previous literature considering insurance and investment problems under constant risk aversion coefficients, when the financial market and the insurance market are independent, the investment strategy is not affected by the parameters of the insurance market and the reinsurance strategy is not influenced by the parameters of the risky asset (see, for instance, the equation (4.44) hereafter). Thus, our newly derived results suggest that taking into account the random risk aversion in reinsurance and investment problems is both meaningful and essential. On the other hand, we find that Li and Li [26] studied a class of mean-variance reinsurance and investment problems under state dependent risk aversion, in which when the risk aversion function exhibits inverse proportionality to wealth, they obtained the semi-analytical equilibrium reinsurance and investment strategy. The expression of their strategy shows that when the financial market and the insurance market are independent, both reinsurance and investment strategies are influenced by parameters of risky assets and insurance market parameters. Although there are similarities between our conclusions and those of [26], we assumed that the risk aversion is a random variable satisfying a certain distribution and derived them under the power utility function by the approach of using expected certainty equivalent.

### v2

@@ 5 Equilibrium strategies under the geometric Brownian motion
In this section, we first attempt to address the delayed reinsurance and investment problem with random risk aversion under the geometric Brownian motion in a general utility function framework as far as possible. Subsequently, in order to obtain (semi-)analytical equilibrium reinsurance and investment strategy, we narrow our focus to analyze the exponential utility function and the power utility function, respectively.
If $\delta=0$, which implies that the CEV model degenerates into geometric Brownian motion, then $U(t,x,s_{1},m_{1})$, $H(t,x,s_{1},m_{1})$ and $y^{\hat{u},\gamma}(t,x,s_{1},m_{1})$ are independent of $s_{1}$. In the sequel, we will adopt the notations $U(t,x,m_{1})$, $H(t,x,m_{1})$ and $y^{\hat{u},\gamma}(t,x,m_{1})$ for simplicity. Then, one can derive the derivatives $H_{t}(t,x,m_{1})$, $H_{x}(t,x,m_{1})$, $H_{m_{1}}(t,x,m_{1})$ and $H_{xx}(t,x,m_{1})$ that are similar to those shown in the equation (4.12). Thus, the equation (3.9) can be rewritten as
$\sup\limits_{u}\;\left\{\mathcal{\widetilde{A}}^{u}U(t,x,m_{1})-0.5(b^{2}q^{2}+\pi^{2}x^{2}\sigma^{2})\int(\iota^{\gamma})^{{}^{\prime}}(Y^{\gamma}(t,x,m_{1}))(Y_{x}^{\gamma}(t,x,m_{1}))^{2}\mathrm{d}\Gamma(\gamma)\right\}=0,$ (5.25)
where
$\mathcal{\widetilde{A}}^{u}\widetilde{\phi}(t,x,m_{1})=$ $\widetilde{\phi}_{t}(t,x,m_{1})+\left[(A+\pi(\mu-r))x+Bm_{1}+Cm_{2}+a\eta+a\eta_{2}q\right]\widetilde{\phi}_{x}(t,x,m_{1})$
$+0.5\left(b^{2}q^{2}+\pi^{2}x^{2}\sigma^{2}\right)\widetilde{\phi}_{xx}(t,x,m_{1})+\left(x-\alpha m_{1}-e^{-\alpha h}m_{2}\right)\widetilde{\phi}_{m_{1}}(t,x,m_{1})$
[…]
@@
Having derived the equation (5.25) that characterizes the equilibrium reinsurance and investment strategy, we can perform partial differentiation on the terms within the brackets of equation (5.25) with respect to $q$ and $\pi$ separately. This step yields the candidate equilibrium reinsurance and investment strategy
@@
$\hat{\pi}(t)=-\frac{(\mu-r)U_{x}(t,x,m_{1})}{\sigma^{2}x\left[U_{xx}(t,x,m_{1})-\int(\iota^{\gamma})^{{}^{\prime}}(Y^{\gamma}(t,x,m_{1}))(Y_{x}^{\gamma}(t,x,m_{1}))^{2}\mathrm{d}\Gamma(\gamma)\right]}.$ (5.27)
It follows from the equation (C) that $U(t,x,m_{1})=H(t,x,m_{1})=\int(\varphi^{\gamma})^{-1}\left(Y^{\gamma}(t,x,m_{1})\right)\mathrm{d}\Gamma(\gamma).$ Then, through the system of equations (4.12), equations (5.26) and (5.27) admit the following reformulation
@@ 5.1 The exponential utility
If $\varphi^{\gamma}$ follows the exponential utility given by the equation (4.17), then equations (5.28) and (5.29) reduce to
@@
We postulate an ansatz where $Y^{\gamma}(t,x,m_{1})=-\frac{1}{\gamma}e^{g_{1}^{\gamma}(t)(x+\beta m_{1})+g_{2}^{\gamma}(t)}$ with $g_{1}^{\gamma}(t)$ and $g_{2}^{\gamma}(t)$ being deterministic functions of $t\in[0,T]$ with $g_{1}^{\gamma}(T)=-\gamma$ and $g_{2}^{\gamma}(T)=0$.
@@
$\left\{\begin{aligned} &Y^{\gamma}_{t}(t,x,m_{1})=Y^{\gamma}(t,x,m_{1})\left[\frac{\partial g_{1}^{\gamma}(t)}{\partial t}\left(x+\beta m_{1}\right)+\frac{\partial g_{2}^{\gamma}(t)}{\partial t}\right],\quad Y^{\gamma}_{x}(t,x,m_{1})=g_{1}^{\gamma}(t)Y^{\gamma}(t,x,m_{1}),\\ &Y^{\gamma}_{xx}(t,x,m_{1})=(g_{1}^{\gamma}(t))^{2}Y^{\gamma}(t,x,m_{1}),\quad Y^{\gamma}_{m_{1}}(t,x,m_{1})=\beta g_{1}^{\gamma}(t)Y^{\gamma}(t,x,m_{1})\end{aligned}\right$.
@@
Inserting the derivatives of $Y^{\gamma}(t,x,m_{1})$ into the equation (3.10), where $\mathcal{A}^{u}$ is replaced by $\mathcal{\widetilde{A}}^{u}$, we have
@@
where $\hat{q}$ and $\hat{\pi}$ are provided by equations (5.30) and (5.31). Then, under the assumption $(A2)$, the above equation arrives at
@@
One can see that $\hat{q}$ and $\hat{\pi}x$ are independent of $x$ and $m_{1}$ due to equations (5.30) and (5.31). Thus, applying the coefficient comparison method to the terms containing $x+\beta m_{1}$ in the equation (5.1) generates
@@
where $D=\gamma\left[\frac{(\mu-r)^{2}}{\sigma^{2}\mathbb{E}[\gamma]}+\frac{a^{2}\eta^{2}_{2}}{b^{2}\mathbb{E}[\gamma]}\right]-0.5\gamma^{2}\left[\frac{(\mu-r)^{2}}{\sigma^{2}\mathbb{E}^{2}[\gamma]}+\frac{a^{2}\eta^{2}_{2}}{b^{2}\mathbb{E}^{2}[\gamma]}\right].$ By using the equation (C), one can further derive that the candidate equilibrium value function is
@@
In addition, provided the expression for the candidate equilibrium reinsurance and investment strategy, one can present the following sensitivity analysis on the effects of the delay parameters on $(\hat{q}(t),\hat{\pi}(t)x)$.
@@
$sgn\left(\frac{\partial\hat{f}(t)}{\partial\alpha}\right)=\begin{cases}1,\quad\quad\text{for}\;\quad\,\alpha>-\frac{1}{h}\ln\frac{1}{h},\\ 0,\quad\quad\text{for}\;\quad\,\alpha=-\frac{1}{h}\ln\frac{1}{h},\\ -1,\quad\;\text{for}\;\quad\,\alpha<-\frac{1}{h}\ln\frac{1}{h},\end{cases}$ (5.38)
@@
$sgn\left(\frac{\partial\hat{f}(t)}{\partial\beta}\right)=\begin{cases}1,\quad\quad\text{for}\;\quad\,h<-\frac{1}{\alpha}\ln(1-r-\alpha),\\ 0,\quad\quad\text{for}\;\quad\,h=-\frac{1}{\alpha}\ln(1-r-\alpha),\\ -1,\quad\;\text{for}\;\quad\,h>-\frac{1}{\alpha}\ln(1-r-\alpha).\end{cases}$ (5.39)
@@
Assume that the utility function $\varphi^{\gamma}$ is given by the equation (4.17) with $\gamma\in[\epsilon_{1},\infty)$ and the assumption $(A2)$ holds. Then the equilibrium reinsurance and investment strategy and the corresponding equilibrium value function are given by equations (5.35) and (5.37), respectively.
@@
Note that the equilibrium reinsurance strategy in (5.41) is consistent with the one in the equation (2.9) of [2] and the equilibrium investment amount $\hat{\pi}(t)x$ in (5.41) coincides with the ones with $\beta=0$ in equations (3.11) and (3.14) of [1], which means that under singular risk aversion in the Black-Scholes financial market, the equilibrium strategy solving the problem (3.5) is mathematically equivalent to the optimal strategy solving the problem (3.4).
@@ 5.2 The power utility
We propose an ansatz in which $Y^{\gamma}(t,x,m_{1})=\frac{1}{1-\gamma}(g^{\gamma}(t))^{\gamma}(x+\beta m_{1})^{1-\gamma}$ with $g^{\gamma}(t)$ being a deterministic function of $t\in[0,T]$ with $g^{\gamma}(T)=1$.
@@
$\left\{\begin{aligned} &Y^{\gamma}_{t}(t,x,m_{1})=\frac{\gamma}{1-\gamma}(g^{\gamma}(t))^{\gamma-1}(x+\beta m_{1})^{1-\gamma}\frac{\partial g^{\gamma}(t)}{\partial t},\quad Y^{\gamma}_{x}(t,x,m_{1})=(g^{\gamma}(t))^{\gamma}(x+\beta m_{1})^{-\gamma},\\ &Y^{\gamma}_{xx}(t,x,m_{1})=-\gamma(g^{\gamma}(t))^{\gamma}(x+\beta m_{1})^{-\gamma-1},\quad Y^{\gamma}_{m_{1}}(t,x,m_{1})=\beta(g^{\gamma}(t))^{\gamma}(x+\beta m_{1})^{-\gamma}.\end{aligned}\right$.
Substituting the above derivatives into the candidate equilibrium reinsurance and investment strategy yields
@@
In addition, replacing $\mathcal{A}^{u}$ by $\mathcal{\widetilde{A}}^{u}$ in the equation (3.10) and inserting the above partial derivatives of $Y^{\gamma}(t,x,m_{1})$ into the equation (3.10) allow us to derive
@@
It is worth mentioning that equations (5.48) and (5.2) demonstrate that because of the linear architecture of the ODE for $g^{\gamma}(t)$, which is presented in the equation (5.46), we are able to transform the infinite-dimensional system of ODEs into a description using a single equation. Nevertheless, the equation (5.2) implicitly defines $\varpi(t,\Gamma)$ through a nonlinear operator equation. While (semi-)closed-form solutions for $\varpi(t,\Gamma)$ may exist under special distributional assumptions on $\Gamma$, the general case resists analytical treatment. Therefore, we attempt to discretize $\Gamma$. For discrete distributions, the outer integral over $\Gamma$ reduces to a finite sum, transforming the system of ODEs into a more tractable form.
@@ 5.2.1 Special case with $n$ possible risk aversions
Suppose the random risk aversion is considered to conform to an $n$-point distribution having possible outcomes $\gamma_{i}>0\;(i=1,\cdot\cdot\cdot,n)$, which occur with probabilities $p_{i}=p(\gamma_{i})$ and $\sum\nolimits_{i=1}^{n}p_{i}=1$. Then, one has
@@
and the corresponding equilibrium value function is $V(t,x,m_{1})=(x+\beta m_{1})\sum\limits_{i=1}^{n}(g^{\gamma_{i}}(t))^{\frac{\gamma_{i}}{1-\gamma_{i}}}p_{i},$ where $\varpi(t,\Gamma)$ is given by the equation (5.50).
@@
If $\alpha=h=B=C=\beta=0$, which indicates that there is no inclusion of any delay within the model, then the equilibrium reinsurance and investment strategy becomes $\left(\hat{q}(t),\hat{\pi}(t)\right)=\left(\frac{a\eta_{2}x}{b^{2}\varpi(t,\Gamma)},\frac{(\mu-r)}{\sigma^{2}\varpi(t,\Gamma)}\right)$ and the corresponding equilibrium value function is characterized by $V(t,x)=x\sum\limits_{i=1}^{n}(g^{\gamma_{i}}(t))^{\frac{\gamma_{i}}{1-\gamma_{i}}}p_{i},$ where $\varpi(t,\Gamma)$ is expressed by the equation (5.50) with $g^{\gamma_{i}}(t)$ satisfying
@@
It follows from equations (5.50) and (5.51) that the solution to $g^{\gamma_{i}}(t)$ is influenced by the model parameters of both the risky asset and the insurance market. As a result, $\varpi(t,\Gamma)$ and $(\hat{q}(t),\hat{\pi}(t))$ are also affected by these parameters. This implies that under the Black-Scholes financial market, when the financial market and the insurance market are independent, the equilibrium investment strategy is influenced not only by the parameters of the risky asset but also by those of the insurance market. The same phenomenon exists for the equilibrium reinsurance strategy. However, these phenomena do not exist in the case of the exponential utility function with random risk aversion (see Theorem 5.1). Furthermore, we would like to mention that in previous literature considering insurance and investment problems under constant risk aversion coefficients, when the financial market and the insurance market are independent, the investment strategy is not affected by the parameters of the insurance market and the reinsurance strategy is not influenced by the parameters of the risky asset (see, for instance, the equation (5.54) hereafter). Thus, our newly derived results suggest that taking into account the random risk aversion in reinsurance and investment problems is both meaningful and essential. On the other hand, we find that Li and Li [26] studied a class of mean-variance reinsurance and investment problems under state dependent risk aversion, in which when the risk aversion function exhibits inverse proportionality to wealth, they obtained the semi-analytical equilibrium reinsurance and investment strategy. The expression of their strategy shows that when the financial market and the insurance market are independent, both reinsurance and investment strategies are influenced by parameters of risky assets and insurance market parameters. Although there are similarities between our conclusions and those of [26], we assumed that the risk aversion is a random variable satisfying a certain distribution and derived them under the power utility function by the approach of using expected certainty equivalent.